OpenStudy (anonymous):

It is estimated that a certain job will take a total of 50 person-hours. If 5 people are assigned to this job, how long will it take them to complete the job?

OpenStudy (anonymous):

\[\frac{50p}{h}\]\[5p\]

OpenStudy (anonymous):

get rid of p by dividng 5p/(50p/h)

OpenStudy (anonymous):

A 10 hours B. 40 hours C. 55 hours D. 250 hours

OpenStudy (anonymous):

\[\frac{50p}{h} \div 5p=\frac{50p}{h}.\frac{1}{5p}=\frac{10}{h}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ten hours